OpenStudy (anonymous):

how do I rewrite the expression log2 x + 5log2 y - 5log2 z... as a single logarithm log2 A...?

OpenStudy (hoblos):

log2(x) + 5log2(y) - 5log2(z) = log2(x) + log2(y^5) - log2(z^5) = log2 [ xy^5 / z^5 ]